OpenStudy (anonymous):

Can you solve and explain the steps you would use in solving two step equations.

OpenStudy (chrisplusian):

Example?

OpenStudy (anonymous):

2x - 3 =11 +3 +3 2x = 14 2x/2 14/2 x = 7

OpenStudy (anonymous):

Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

@geny55

OpenStudy (anonymous):

do you understand?

OpenStudy (chrisplusian):

It seems that you have it correct.

OpenStudy (chrisplusian):

Are you asking how do you know what to do first?

OpenStudy (chrisplusian):

If so the solution to these problems comes from the idea that an equal sign allows you to do anything to one side of an equation as long as you do the same thing to the other side.... defining a side means the whole side of the equation not just part of it. However some things make solving these equations easier than others example.....

OpenStudy (chrisplusian):

2x-3=11

OpenStudy (chrisplusian):

\[\frac{ 2x-3 }{ 2 }=\frac{ 11 }{ 2 }\]

OpenStudy (chrisplusian):

\[x-\frac{ 3 }{ 2 }=\frac{ 11 }{ 2 }\]

OpenStudy (chrisplusian):

\[x-\frac{ 3 }{ 2 }+\frac{ 3 }{ 2 }=\frac{ 11 }{ 2 }+\frac{ 3 }{ 2 }\]

OpenStudy (chrisplusian):

\[x=\frac{ 14 }{ 2 }\]

OpenStudy (chrisplusian):

\[x=7\]

OpenStudy (chrisplusian):

Same answer but what I did was much more difficult than what you did. Typically try to move all constant terms (numbers that aren't attached to a variable) to one side and all variable terms to the other. Then do division or multiplication as needed

OpenStudy (chrisplusian):

Does that make sense? If ever you find yourself in a situation where you have complicated things like I did above, then simply stick to the rule that IF your dealing with an "equation" (meaning there is an equal sign somewhere in your problem) then you can do anything.... anything to one entire side, as long as you do the same thing to the entire other side

OpenStudy (chrisplusian):

Hope that helps

OpenStudy (anonymous):

so can you give me the steps like this :) 1.... 2... 3....

OpenStudy (anonymous):

and explain the steps in words please :)